OpenStudy (anonymous):

Use the elimination method to solve the system. 2x – 2y = –8 x + 2y = –1

OpenStudy (anonymous):

@uri

OpenStudy (uri):

Add both the equations.

OpenStudy (uri):

Remember -2y and +2y will be cancelled.

OpenStudy (anonymous):

kk

OpenStudy (uri):

@Ashley_Cute add and show,what answer you get?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

okay i got 3x-9

OpenStudy (uri):

3x=-9*

OpenStudy (anonymous):

o sry

OpenStudy (uri):

Now divide -9 with 3,What do you get?

OpenStudy (anonymous):

-3

OpenStudy (uri):

Correct,So x is -3 Now put the value of x in this eq: x + 2y = –1

OpenStudy (uri):

^second eq

OpenStudy (anonymous):

kk

OpenStudy (uri):

We have to find y now

OpenStudy (anonymous):

would i subtract 3 or add 1

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

2y=-1+ 3 :)

OpenStudy (anonymous):

@Ashley_Cute

OpenStudy (anonymous):

kk so in order to get y by itself i add 1

OpenStudy (uri):

x + 2y = –1 -3 +2y=-1 We will shift -3 to the right side so sign changes 2y=-1+3 2y=2 Now divide 2/2 and answer

OpenStudy (anonymous):

y=1 :)

OpenStudy (anonymous):

thx guys

OpenStudy (uri):

Yw :)

OpenStudy (anonymous):

wc